The aggregate wage bill in an economy is equal to 60 and output, which is produced according to the Cobb-Douglas production function, is equal to 100. The output growth rate was 10 percent and the growth rates of capital and labor were 10 percent and 5 percent respectively. A) What was the overall productivity (TFP) growth rate for this economy? B) Repeat (a) if the wage bill is 80 instead of 60. C) Derive the expression for TFP growth in an economy that produces output according Y= (AK)NTi-o-p, where T denotes a fixed to the following production function: amount of arable land.
Answer #1

Answer:

OL = 10一千一5-3% thus TFf ghouwth Rate 3% du 8) Now wage Biul 8, then (e)- 心。2 = 10-2-4 = 490 dav C) Now y:@sy( N β.(T) 1-0(-β , Tfr gsoeotafoo TPP = Multiply 8harreof apittoroltoth Rate af technobg